# Pritam Goswami    Adri Bhattacharya    Raja Das    Partha Sarathi Mandal

### v1

@@ 5.2 Communication: Whiteboard
In this section, we discuss the algorithm PerpExplore-Scat-WhitBrd, which achieves PerpExploration-BBH, with the help of three agents on a ring $R$ with $n$ nodes, where each node has a whiteboard that can store $O(\log n)$ bits of data. Note that, in this model as well we have considered that the agents are placed arbitrarily placed along the nodes of the ring $R$ (each such node is a ‘safe node’). So, there can be two cases, first, all the three agents are initially placed at three different positions, second, two agents are together whereas the third agent is in a different position. We will provide algorithm for the first case only, as an algorithm for the second case can be easily designed by modifying and merging the algorithms for the first case and algorithm PerpExplore-Coloc-Pbl.
The idea of our algorithm PerpExplore-Scat-WhitBrd is as follows: the three agents separately explore three segments of the ring $R$, such that, their union is the whole ring $R$ and intersection of any two segment contains only a safe node (which is a $home$ for one agent). Thus, while exploring, only one agent can be consumed by the black hole, and in an additional finite number of rounds another agent can learn,the consumption of the earlier agent by the black hole (this agent is precisely the agent whose segment has a safe node common with the segment of the consumed agent. Also it learns about this consumption from the whiteboard of the common node). In which case, this agent gathers with the other alive agent, and locates the exact segment in which the earlier agent was exploring before its consumption by the black hole. After locating this segment, both these agents starts performing Cautious-Walk either in clockwise or counter-clockwise direction, based on the information they gained about the earlier agents movement direction before getting consumed by the black hole. The agents before getting consumed, marks each node of their corresponding segment with different markings (i.e., more precisely writes either left or right, after erasing the earlier direction marking at the current node) based on the direction of their movement. So, the two agents performing Cautious-Walk, search for the first node, which does not have the direction marking w.r.to the direction the earlier consumed agent was moving. This node which is devoid of this direction marking is indeed the black hole node, and at least one agent, among the agents performing Cautious-Walk detects it, and remains alive, which further perpetually explores the ring $R$, avoiding this node.
@@
Let us consider the three agents (say, $a_{1}$, $a_{2}$ and $a_{3}$) are initially placed at three nodes of the ring $R$, which are not only safe nodes but are also recognised as the $home$ of these agents. Initially, the agents starting from their respective $home$ are assigned the task to explore a set of nodes, which we term as a segment of the corresponding agents. More precisely, a segment for an agent $a_{i}$ is defined as the set of consecutive nodes, starting from its $home$ and ending at the nearest clockwise $home$ (i.e., the $home$ of first clockwise placed agent), which is also termed as $Seg(a_{i})$. Note that $\cup^{3}_{i=1}Seg(a_{i})=R$. So if none of the agents are ever destroyed the ring will still be perpetually explored.
@@
Let us first discuss the case, in which we describe the possible movements of the agent and the respective state changes they perform, until one agent gets consumed by the black hole, and another agent gets to know that an agent is already consumed by the black hole. After which, we will describe all the possible movements and state changes performed by the remaining two alive agent, between getting to know that already one has been consumed by the black hole, and finally detecting the black hole position.
@@
Initially all agents start from state Initial at their respective $home$. In state Initial, an agent ( without loss of generality, say $a_{1}$) first clears the already present data (if at all) at the whiteboards of their respective $home$, then initializes $T_{time}=0$ ($T_{time}$ stores the number of rounds elapsed since the start of state Initial), and writes a message of the form (home, ID) at its $home$, where ID is the ID of the agent. This type of message is termed as “home" type message, which consists of two components, first component stores the message home and second component stores the ID of the agent writing, i.e., ID of the agent whose $home$ is the $Current-Node$. Further, it changes its state to Forward, and moves in clockwise direction.
@@
In state Forward, the agent moves in clockwise direction while erasing the earlier direction marking (if exists), i.e., left and then writes the new direction marking, i.e., right in each node. The agent also increases the $T_{time}$ variable by one in each round. This process continues until the agent encounters a node which has a “home" type message. This “home" type message signifies that the agent has reached the end of segment $Seg(a_{1})$, i.e., in other words it has reached the nearest clockwise $home$, say $v_{c}$. Note that the length of a segment can be at most $n-2$, hence within $T_{time}=n-1$, an agent $a_{1}$ is bound to reach the last node of its own segment i.e., $v_{c}$. In any case irrespective of the current $T_{time}$, the agent waits at $v_{c}$ until $T_{time}=n-1$, after which in the next round, it checks for the following information at $v_{c}$. If the agent finds a message of type “visited" at $v_{c}$, the agent considers this as an anomaly and learns that an agent of which $v_{c}$ is the home (from the ID component of home type message at $v_{c}$) must have entered black hole while returning back from its clockwise nearest $home$ (refer Lemma 5.9). Then in this case, the agent stores the message of type dir in its local memory, where dir=(Counter-clockwise, NULL). In a dir type message, the first component is called a direction component which indicates the direction of the agent which gets consumed by the black hole, along which it was moving just before it gets consumed. On the other hand, the second component either stores ID of some agent or it stores $NULL$ message. Both these components are useful for certain state transitions. Next, the agent changes its state to Gather. Otherwise, if no anomaly is detected, then the agent simply writes the message (Visited, self ID) at $v_{c}$ and changes its state to Back-Wait.
@@
Next, in state Back-Wait, an agent (here $a_{1}$) waits at $v_{c}$ until $T_{time}=2n-1$. In this waiting time $a_{1}$ waits for other agent to meet $a_{1}$ if the other agent detects any anomaly during its state Forward. While waiting, if it finds that the #$\#Agent$ at $v_{c}$ is more than 1, and the whiteboard at $v_{c}$ has a message of type dir, then the agent performs the following task. After noticing this, the agent moves along the counter clockwise direction, while changing its state to Gather2. Note that, our algorithm ensures that in state Back-Wait, if a dir type message is seen by an agent, then the direction component of this message must be in a counter-clockwise direction (refer Lemma 5.15). Otherwise, if no such dir type message is seen by $a_{1}$, and also $T_{time}=2n$, then in this round, the agent changes its state to Backtrack.
@@
In state Backtrack, an agent (here $a_{1}$) starts to move in a counter-clockwise direction from $v_{c}$, and after each move erases the earlier direction, i.e., right, and writes the new direction, i.e., left, and also increments $T_{time}$ by 1. This process continues until it reaches its own $home$, i.e., reads a home type message with ID same to its own ID. Again note that, since a segment can be of length at most $n-2$, hence within $T_{time}=3n-1$, the agent reaches its own $home$. After which it does not move until $T_{time}=3n$. At $T_{time}=3n$, it checks whether its $home$ has visited type message, if so then directly changes its state to Initial-Wait. Otherwise, absence of such visited type message, creates an anomaly for the agent. This only happens when another agent, say $a_{j}$, ($j\neq 1$), for which $Seg(a_{1})\cap Seg(a_{j})=$ $home$ of $a_{1}$, does not arrive at $home$ of $a_{1}$ during its Forward state because of getting consumed at the black hole (refer Lemma 5.11). This instigates the agent $a_{1}$ to change its state to Gather and store the dir type message (clockwise,NULL) in its local memory..
@@
In state Initial-Wait, the agent waits at its $home$ until $T_{time}=4n-1$. This waiting period is enough for an anomaly finding agent in state Backtrack to meet with it. If the agent finds that there are more than one agent at its $home$, and also there is a dir type message as well, then our algorithm ensures that the direction component of this dir type message must be in clockwise direction (refer Lemma 5.17). In which case, the agent starts moving clockwise and changes its state to Gather1. On the contrary, if $T_{time}=4n$ and no anomaly is detected, then the agent again moves back to state Initial.
@@
An agent can change its state to Gather, either from state Forward, or from state Backtrack. Note that in either case, it carries the stored message of type dir (which has a direction component and an ID component which is either $NULL$ or stores the ID of an agent). So, in this state if the direction component of dir is along clockwise (resp. counter-clockwise), then the agent moves along clockwise (resp. counter-clockwise) direction until it finds a node with #$\#Agent$ more than one. Next, in case of counter-clockwise direction component, the agent updates the dir message to (Counter-clockwise, ID’) where ID’ is the ID of the other agent at the same node. Then the agent writes the updated dir message at the current node and changes state to Gather2. If the agent in state Gather met the other agent when the dir type message having clockwise direction component the, the agent simply just writes the message at the current node and moves to state Gather1.
@@
Also note that exactly one agent can move into state Gather. This is because if an agent is consumed during Forward state then the anomaly is first found by a single agent whose $home$ is the clockwise nearest $home$ of the consumed agent. This agent changes its state to Gather. Now, before the other agent finds further anomaly the agent in state Gather meets the other agent and forces it to change into state Gather1. Similarly If an agent $a_{1}$ is consumed by the black hole during its Backtrack state, then the first anomaly is detected by a single agent (more precisely the agent for which the clockwise nearest $home$ is the $home$ of $a_{1}$). In this case it changes its state to Gather and forces the other agent to move into state Gather2 by meeting it before it finds further anomaly. Thus we have the following observation. {observation} There is exactly one agent that changes its state to Gather throughout the execution of Algorithm PerpExplore-Scat-Whitbrd. Also, The agent which changes its state to Gather, is the first agent to understand that, an agent has already been consumed by the black hole.
@@
An agent can change its state to Gather1, only from the states Initial-Wait or Gather. An agent changes its state to Gather1 from Gather, only when while moving along clockwise direction finds another agent present, in which case it writes the corresponding message of type dir and correspondingly changes its state to Gather1. on the other hand, an agent (here $a_{1}$) can only change its state to Gather1 from Initial-Wait, if it is waiting at its $home$, and within which it encounters another agent at its $home$ along with it a message of type dir is written (in which the direction component of dir is along clockwise). Note that if $a_{1}$ in state Initial-Wait finds more than one agent at current node and dir type message at round $t$ then, the other agent must be in state Gather while written there at round $t-1$ and moved into state Gather1 in the same round. Then, during round $t$, $a_{1}$ first stores the dir type message before it moves to state Gather1 and moves according to the direction component of the dir type message (i.e., clockwise). Also, the other agent that changed its state to Gather1 from Gather in round $t-1$ also moves according to the dir type message during round $t$. Thus we have another observation {observation} If two agents move to state Gather1 , one moves from state Gather and the other moves from state Initial-Wait. Furthermore, they change to state Gather1 at the same node and leaves the node together at the same round and stays together.
@@
In state Gather1, the agent which changes its state to Gather1 from Initial-Wait first stores the dir type message, as the other agent must already have stored this message in state Gather. Irrespective of which, the agent starts moving in a clockwise direction, until it finds a node which has a home (i.e., of the form (home, ID)) type message, where the ID component this message does not match with the IDs of the agents present at the current node. Note that our algorithm ensures that the current node is the $home$ of $a_{j}$, if $a_{j}$ is the agent to be consumed by the black hole (refer Lemma 5.19). Further, as there are two agents at the current node both in state Gather1 (Observation 5.2), so now the agent with the lowest ID, sets $Move=0$, also initializes the $Marking$ variable to $right$, since the direction component of dir is clockwise, and further changes its state to Cautious-Leader. On the other hand the other alive agent at the current node changes its state to Cautious-Follower, while initializing $Move=0$.
An agent can change its state to Gather2, only from the states Back-Wait or Gather.
An agent changes its state to Gather2 from Gather, only when while moving along counter-clockwise direction finds another agent present, in which case it updates the dir message and writes the corresponding message of type dir and correspondingly changes its state to Gather2.
On the other hand, an agent can only change its state to Gather2 from Back-Wait, if it is waiting at its corresponding segments nearest clockwise $home$, and within which it encounters another agent at its current node along with it a message of type dir is written (in which the direction component of dir is along counter-clockwise). After storing the dir type message in its local memory and changing its state to Gather2 from Back-Wait, an agent immediately moves counter-clockwise in the same round. This generates another observation similar to the Observation 5.2 as follows.
{observation}
If two agents move to state Gather2 , one moves from state Gather and the other moves from state Back-Wait. Furthermore, they change to state Gather2 at the same node and leaves the node together at the same round and stays together.
In state Gather2, the agent moves in a counter-clockwise direction, until it finds a node which has a home (i.e., of the form (home, ID’)) type message, where ID’ matches with ID component of dir type message in its local memory. Note that our algorithm ensures that the current node is the nearest clockwise $home$ of $a_{1}$, if $a_{1}$ is the agent to be consumed by the black hole (refer Lemma 5.19). Further, as there are two agents at the current node, both in state Gather2 (refer Observation 5.2), so now the agent with the lowest ID, sets $Move=0$, also initializes the $Marking$ variable to $left$, since the direction component of dir is counter-clockwise, and further changes its state to Cautious-Leader. On the other hand the other alive agent at the current node, changes its state to Cautious-Follower, while initializing $Move=0$.
In state Cautious-Leader, an agent updates $Move=1$ and moves along the direction component of dir. After moving when it is alone on a node, it checks if the current node is marked with $Marking$ (i.e., either $left$ or $right$). If the node is marked, then it moves in the opposite direction (i.e., the opposite of direction component of dir message) to meet with the agent in state Cautious-Follower. Otherwise, if there is no $Marking$ at the current node then it identifies the current node as the black hole node (only if it remains alive) and continues perpetual exploration avoiding this node. If an agent in state Cautious-Leader with $Move=1$ meets with another agent (i.e., the agent in state Cautious-Follower) it updates variable $Move$ to 0, and again moves according to the direction component.
In state Cautious-Follower, the agent after finding $Move=0$, initializes the variable $Wait=0$ and updates $Move=1$. If $Move=1$, then the agent checks whether $Wait<1$, if so then update $Wait=Wait+1$. When $Wait=1$, it checks if the current node has more than one agent (i.e., whether the agent in state Cautious-Leader has returned or not), if so then further update $Move=0$ and move along the direction component of dir, along with the agent in state Cautious-Leader. Otherwise, if the agent in state Cautious-Leader does not return, that means the number of agent at the current node is 1 when $Wait=1$, this symbolizes that the agent in state Cautious-Leader has been either consumed by the black hole or the agent in state Cautious-Leader finds that its current node (i.e., the next node along the direction component of dir for the agent in state Cautious-Follower) is the black hole and continues to perpetually move in the same direction, further avoiding this node. In any case, the agent executing Cautious-Leader has not returned implies the next node along direction component of dir for the agent executing Cautious-Follower is the black hole node. So, in this case, the agent declares the next node along the direction component of dir is the black hole node and continues perpetual exploration avoiding that node.

### v2

@@ 11 Algorithm PerpExplore-Scat-Whitbrd
In this section, we discuss the algorithm PerpExplore-Scat-WhitBrd, which achieves PerpExploration-BBH, with the help of three agents on a ring $R$ with $n$ nodes, where each node has a whiteboard that can store $O(\log n)$ bits of data. Note that, in this model as well we have considered that the agents are placed arbitrarily placed along the nodes of the ring $R$ (each such node is a ‘safe node’). So, there can be two cases, first, all the three agents are initially placed at three different positions, second, two agents are together whereas the third agent is in a different position. We will provide algorithm for the first case only, as an algorithm for the second case can be easily designed by modifying and merging the algorithms for the first case and algorithm PerpExplore-Coloc-Pbl (as discussed in Remark 11.15 in Appendix, Section 11.3).
11.1 Detailed description and Pseudocode of the algorithm
@@
Let us consider the three agents (say, $a_{1}$, $a_{2}$ and $a_{3}$) are initially placed at three nodes of the ring $R$, which are not only safe nodes but are also recognized as the $home$ of these agents. Initially, the agents starting from their respective $home$ are assigned the task to explore a set of nodes, which we term as a segment of the corresponding agents. More precisely, a segment for an agent $a_{i}$ is defined as the set of consecutive nodes, starting from its $home$ and ending at the nearest clockwise $home$ (i.e., the $home$ of first clockwise placed agent), which is also termed as $Seg(a_{i})$. Note that $\cup^{3}_{i=1}Seg(a_{i})=R$. So if none of the agents are ever destroyed the ring will still be perpetually explored.
@@
Let us first discuss the case, in which we describe the possible movements of the agent and the respective state changes they perform, until one agent gets destroyed by the Byzantine black hole, and another agent gets to know that an agent is already destroyed by the Byzantine black hole. After which, we will describe all the possible movements and state changes performed by the remaining two alive agents, between getting to know that already one has been destroyed by the Byzantine black hole, and finally detecting the Byzantine black hole’s position.
@@
Initially, all agents start from state Initial at their respective $home$. In state Initial, an agent ( without loss of generality, say $a_{1}$) first clear the already present data (if at all) at the whiteboards of their respective $home$, then initializes $T_{time}=0$ ($T_{time}$ stores the number of rounds elapsed since the start of state Initial), and writes a message of the form (home, ID) at its $home$, where ID is the ID of the agent. This type of message is termed as “home" type message, which consists of two components, the first component stores the message home and the second component stores the ID of the agent writing, i.e., ID of the agent whose $home$ is the $Current-Node$. Further, it changes its state to Forward and moves in the clockwise direction.
@@
In state Forward, the agent moves in the clockwise direction while erasing the earlier direction marking (if exists), i.e., left and then writes the new direction marking, i.e., right in each node. The agent also increases the $T_{time}$ variable by one in each round. This process continues until the agent encounters a node that has a “home" type message. This “home" type message signifies that the agent has reached the end of segment $Seg(a_{1})$, i.e., in other words, it has reached the nearest clockwise $home$, say $v_{c}$. Note that the length of a segment can be at most $n-2$, hence within $T_{time}=n-1$, an agent $a_{1}$ is bound to reach the last node of its own segment i.e., $v_{c}$. In any case irrespective of the current $T_{time}$, the agent waits at $v_{c}$ until $T_{time}=n-1$, after which in the next round, it checks for the following information at $v_{c}$. If the agent finds a message of type “visited" at $v_{c}$, the agent considers this as an anomaly and learns that an agent of which $v_{c}$ is the home (from the ID component of home type message at $v_{c}$) must have entered Byzantine black hole while returning back from its clockwise nearest $home$ (refer Lemma 11.1). Then in this case, the agent stores the message of type dir in its local memory, where dir=(Counter-clockwise, NULL). In a dir type message, the first component is called a direction component which indicates the direction of the agent that gets destroyed by the Byzantine black hole, along which it was moving just before it gets destroyed. On the other hand, the second component either stores ID of some agent or stores $NULL$ message. Both these components are useful for certain state transitions. Next, the agent changes its state to Gather. Otherwise, if no anomaly is detected, then the agent simply writes the message (Visited, self ID) at $v_{c}$ and changes its state to Back-Wait.
@@
Next, in state Back-Wait, an agent (here $a_{1}$) waits at $v_{c}$ until $T_{time}=2n-1$. In this waiting time, $a_{1}$ waits for the other agent to meet $a_{1}$ if the other agent detects any anomaly during its state Forward. While waiting, if it finds that the #$\#Agent$ at $v_{c}$ is more than 1, and the whiteboard at $v_{c}$ has a message of type dir, then the agent performs the following task. After noticing this, the agent moves along the counter-clockwise direction, while changing its state to Gather2. Note that, our algorithm ensures that while in state Back-Wait, if a dir type message is seen by an agent, then the direction component of this message must be in a counter-clockwise direction (refer Lemma 11.7). Otherwise, if no such dir type message is seen by $a_{1}$, and also $T_{time}=2n$, then in this round, the agent changes its state to Backtrack.
@@
In state Backtrack, an agent (here $a_{1}$) starts to move in a counter-clockwise direction from $v_{c}$, and after each move erases the earlier direction, i.e., right, and writes the new direction, i.e., left, and also increments $T_{time}$ by 1. This process continues until it reaches its own $home$, i.e., reads a home type message with ID same as its own ID. Again note that, since a segment can be of length at most $n-2$, hence within $T_{time}=3n-1$, the agent reaches its own $home$. After which it does not move until $T_{time}=3n$. At $T_{time}=3n$, it checks whether its $home$ has visited type message, if so then directly changes its state to Initial-Wait. Otherwise, the absence of such visited type message, creates an anomaly for the agent. This only happens when another agent, say $a_{j}$, ($j\neq 1$), for which $Seg(a_{1})\cap Seg(a_{j})=$ $home$ of $a_{1}$, does not arrive at $home$ of $a_{1}$ during its Forward state because of getting destroyed at the Byzantine black hole (refer Lemma 11.3). This instigates the agent $a_{1}$ to change its state to Gather and store the dir type message (clockwise, NULL) in its local memory.
@@
In state Initial-Wait, the agent waits at its $home$ until $T_{time}=4n-1$. This waiting period is enough for an anomaly finding agent in state Backtrack, to meet with it. If the agent finds that there is more than one agent at its $home$, and also there is a dir type message as well, then our algorithm ensures that the direction component of this dir type message must be in clockwise direction (refer Lemma 11.9). In this case, the agent starts moving clockwise and changes its state to Gather1. On the contrary, if $T_{time}=4n$ and no anomaly is detected, then the agent again moves back to state Initial.
@@
An agent can change its state to Gather, either from state Forward or from state Backtrack. Note that in either case, it carries the stored message of type dir (which has a direction component and an ID component that is either $NULL$ or stores the ID of an agent). So, in this state if the direction component of dir is along clockwise (resp. counter-clockwise), then the agent moves along clockwise (resp. counter-clockwise) direction until it finds a node with #$\#Agent$ more than one. Next, in case of a counter-clockwise direction component, the agent updates the dir message to (Counter-clockwise, ID’) where ID’ is the ID of the other agent at the same node. Then the agent writes the updated dir message at the current node and changes state to Gather2. If the agent in the state Gather met the other agent when the dir type message has a clockwise direction component, the agent simply just writes the message at the current node and moves to state Gather1.
@@
Also, note that exactly one agent can move into state Gather. This is because if an agent is destroyed during Forward state then the anomaly is first found by a single agent whose $home$ is the clockwise nearest $home$ of the destroyed agent. This agent changes its state to Gather. Now, before the other agent finds further anomaly the agent in state Gather meets the other agent and forces it to change into state Gather1. Similarly, If an agent $a_{1}$ is destroyed by the Byzantine black hole during its Backtrack state, then the first anomaly is detected by a single agent (more precisely the agent for which the clockwise nearest $home$ is the $home$ of $a_{1}$). In this case, it changes its state to Gather and forces the other agent to move into state Gather2 by meeting it before it finds further anomaly. Thus we have the following observation. {observation} There is exactly one agent that changes its state to Gather throughout the execution of Algorithm PerpExplore-Scat-Whitbrd. Also, The agent which changes its state to Gather, is the first agent to understand that, an agent has already been destroyed by the Byzantine black hole.
@@
An agent can change its state to Gather1, only from the states Initial-Wait or Gather. An agent changes its state to Gather1 from Gather, only when while moving along clockwise direction finds another agent present, in which case it writes the corresponding message of type dir and correspondingly changes its state to Gather1. on the other hand, an agent (here $a_{1}$) can only change its state to Gather1 from Initial-Wait, if it is waiting at its $home$, and within which it encounters another agent at its $home$ along with it a message of type dir is written (in which the direction component of dir is along clockwise). Note that if $a_{1}$ in state Initial-Wait finds more than one agent at the current node and dir type message at round $t$ then, the other agent must be in state Gather while written there at round $t-1$ and moved into state Gather1 in the same round. Then, during round $t$, $a_{1}$ first stores the dir type message before it moves to state Gather1 and moves according to the direction component of the dir type message (i.e., clockwise). Also, the other agent that changed its state to Gather1 from Gather in round $t-1$ also moves according to the dir type message during round $t$. Thus we have another observation {observation} If two agents move to state Gather1, one moves from state Gather, and the other moves from state Initial-Wait. Furthermore, they change to state Gather1 at the same node and leave the node together at the same round and stay together.
@@
In state Gather1, the agent which changes its state to Gather1 from Initial-Wait first stores the dir type message, as the other agent must already have stored this message in state Gather. Irrespective of which, the agent starts moving in a clockwise direction until it finds a node that has a home (i.e., of the form (home, ID)) type message, where the ID component this message does not match with the IDs of the agents present at the current node. Note that our algorithm ensures that the current node is the $home$ of $a_{j}$ if $a_{j}$ is the agent to be destroyed by the Byzantine black hole (refer Lemma 11.11). Further, as there are two agents at the current node both in state Gather1 (Observation 11.1), so now the agent with the lowest ID, sets $Move=0$, also initializes the $Marking$ variable to $right$, since the direction component of dir is clockwise, and further changes its state to Cautious-Leader. On the other hand, the other alive agent at the current node changes its state to Cautious-Follower, while initializing $Move=0$.